# OpenSIR: Open-Ended Self-Improving Reasoner

### v1

@@ Abstract
Recent advances in large language model (LLM) reasoning through reinforcement learning rely on annotated datasets for verifiable rewards, which may limit models’ ability to surpass human-level performance. While self-play offers a promising alternative, existing approaches depend on external verifiers or cannot learn open-endedly. We present Open-Ended Self-Improving Reasoner (OpenSIR), a self-play framework where an LLM learns to generate and solve novel problems by alternating teacher and student roles without external supervision. To generate novel problems, OpenSIR optimises for both difficulty and diversity, rewarding problems that challenge appropriately while exploring distinct concepts, enabling open-ended mathematical discovery. Starting from a single trivial seed problem, OpenSIR substantially improves instruction models: Llama-3.2-3B-Instruct advances from 73.9 to 78.3 on GSM8K, and from 28.8 to 34.4 on College Math, while Gemma-2-2B-Instruct rises from 38.5 to 58.7 on GSM8K. Our analyses reveal that OpenSIR achieves open-ended learning through co-evolving teacher-student roles that adaptively calibrate difficulty and drive diverse exploration, progressing autonomously from basic to advanced mathematics. The code is publicly available at https://github.com/EdinburghNLP/OpenSIR.
@@ 1 Introduction
Reinforcement learning with verifiable rewards (RLVR) drives recent advances in LLM reasoning. Recent works on DeepSeek-R1 (DeepSeek-AI et al., 2025) and OpenAI o1 (OpenAI, 2024) have shown that large-scale reinforcement learning improves reasoning capabilities. Yet these methods need extensive human-annotated data for reward signals, bottlenecking scalability and potentially limiting performance to human-level (Hughes et al., 2024b).
@@ 3.1 Training Setup
We experiment with three instruction-tuned models: Llama-3.2-3B-Instruct (Dubey et al., 2024), Gemma-2-2B-Instruct (Team et al., 2024), and Qwen-2.5-3B-Instruct (Team, 2024) with GRPO (Shao et al., 2024). We use a learning rate of $3\times 10^{-7}$ and 10 warm-up steps. The KL divergence coefficient is set to $10^{-4}$ and the batch size is 256. To compare models trained on the same number of problem-solution pairs, we train the GRPO baselines with 100 steps, and OpenSIR for 200 steps since OpenSIR allocates half of its training budget to problem generation. Clipping is not applied since we strictly use on-policy samples. Each experiment is run with three random seeds. We provide full training details in Appendix D.1.
@@ 3.4 Main Results
OpenSIR 85.38 65.87 25.96 48.74 28.33 50.85+0.6
Table 1: The avg@16 performance on five mathematical benchmarks. OpenSIR consistently outperforms GRPO baselines across model architectures despite generating training data through self-play from a single seed problem, while GRPO baselines use over 7,000 human-annotated examples.
Table 1 demonstrates that OpenSIR achieves substantial gains over the base instruction models, improving Llama-3.2-3B-Instruct by 3.6 points and Gemma-2-2B-Instruct by 5.9 points on average accuracy. While OpenSIR demonstrates huge improvements with Llama-3.2-3B-Instruct and Gemma-2-9b-Instruct, Qwen-2.5-3B-Instruct demonstrates a relatively minimal gains (+0.3). Notably, GRPO baselines exhibit similarly marginal improvements with this model, indicating the effect is not specific to OpenSIR. The limited improvement aligns with observations of potential benchmark contamination (Wu et al., 2025).
OpenSIR outperforms all GRPO baselines without using human-annotated training data. GRPO baselines require over 7,000 labeled examples, yet OpenSIR generates its own training problems through self-play, starting from a single trivial seed problem. As we reveal later in our analysis, the success of OpenSIR can be attributed to its ability to explore diverse mathematical concepts, and calibrating difficulty adaptively to maintain optimal challenge levels (§4.1). These capabilities enable OpenSIR to self-improve and expand its skills without external training data, achieving open-ended learning.
@@ A.1 Full Results
Table 5: Math reasoning evaluation results with individual seed reporting. We report avg@16 per problem for each seed (s-42, s-43, s-44) and their average with standard deviation as superscript.
We provide the full results of all seeds in Table 5.

### v2

@@ Abstract
Recent advances in large language model (LLM) reasoning through reinforcement learning rely on annotated datasets for verifiable rewards, which may limit models’ ability to surpass human-level performance. While self-play offers a promising alternative, existing approaches depend on external verifiers or cannot learn open-endedly. We present Open-Ended Self-Improving Reasoner (OpenSIR), a self-play framework where an LLM learns to generate and solve novel problems by alternating teacher and student roles without external supervision. To generate novel problems, OpenSIR optimises for both difficulty and diversity, rewarding problems that challenge appropriately while exploring distinct concepts, enabling open-ended mathematical discovery. Starting from a single trivial seed problem, OpenSIR substantially improves instruction models: Llama-3.2-3B-Instruct advances from 73.9 to 78.3 on GSM8K, and from 28.8 to 34.4 on College Math, while Gemma-2-2B-Instruct rises from 38.5 to 58.7 on GSM8K. Our analyses reveal that OpenSIR achieves open-ended learning through co-evolving teacher-student roles that adaptively calibrate difficulty and drive diverse exploration, progressing autonomously from basic to advanced mathematics. Code and datasets are available at https://github.com/EdinburghNLP/OpenSIR.
@@ 1 Introduction
Reinforcement learning with verifiable rewards (RLVR) drives recent advances in LLM reasoning. Recent works on DeepSeek-R1 (DeepSeek-AI et al., 2025) and OpenAI o1 (OpenAI, 2024) have shown that large-scale reinforcement learning improves reasoning capabilities. Yet, these methods require extensive human-annotated data for reward signals, which bottleneck scalability and potentially limit performance to human-level (Hughes et al., 2024b).
@@ 2.4 Model Update
Algorithm 1 OpenSIR
1:Problem pool $\mathcal{P}_{0}$, policy $\pi_{\theta}^{(0)}$, embedding encoder $\varepsilon$, batch size $B$, generation group size $G$, solve rate range $s_{\min},s_{\max}]$, teacher prompt $p_{T}$, student prompt $p_{S}$
2:for $t=1$ to $T$ do
3:  Sample $k=B/G$ reference problems $\{p_{1},\ldots,p_{k}\}$ from $\mathcal{P}_{t-1}$ $\triangleright$ Problem Generation
4:  for $i=1$ to $k$ do
5:   Sample $q_{i,1:G}\sim\pi_{\theta}^{(t)}(\cdot\mid p_{i},p_{T})$
6:  end for
7:  $\mathcal{Q}_{\text{valid}}\leftarrow\{q_{i,j}\mid q_{i,j}\text{ has valid format}\}$
8:  for each $q_{i}\in\mathcal{Q}_{\text{valid}}$ do $\triangleright$ Solution Sampling
9:   Sample solutions $o_{i,1:G}\sim\pi_{\theta}^{(t)}(\cdot\mid q_{i},p_{S})$
10:   Parse answers $a_{i,1:G}$ from solutions $o_{i,1:G}$
11:   Compute reference answer $a_{i}^{*}=\arg\max_{a\in a_{i,1:G}}\text{count}(a)$ via majority voting
12:   Compute solve rate $s_{q_{i}}=\text{count}(a_{i}^{*})/G$
13:   Compute embedding $e_{q_{i}}\leftarrow\varepsilon(q_{i})$
14:  end for
15:  Compute $\text{score}_{\text{novel}}(q_{i})$ for all $q_{i}\in\mathcal{Q}_{\text{valid}}$ via Eq. 5 $\triangleright$ Scoring
16:  $\mathcal{I}_{T}\leftarrow\text{top}_{B/(2G)}(i:\text{Var}(\text{score}_{\text{novel}}(q_{i,1:G})),i\in\{1,\ldots,k\})$ $\triangleright$ Teacher sample selection
17:  $\mathcal{Q}_{S}\leftarrow\text{top}_{B/(2G)}(q:\text{score}_{\text{novel}}(q),q\in\mathcal{Q}_{\text{valid}})$ $\triangleright$ Student sample selection
18:  Compute $\text{score}_{\text{correct}}(o_{i,j},a_{i,j})$ for solutions where $q_{i}\in\mathcal{Q}_{S}$ via Eq. 6
19:  $\mathcal{D}_{T}\leftarrow\{(p_{T},q_{i,j},R_{i,j}^{T}):i\in\mathcal{I}_{T},1\leq j\leq G\}$ where $R_{i,j}^{T}=\text{score}_{\text{novel}}(q_{i,j})$ $\triangleright$ Model Update
20:  $\mathcal{D}_{S}\leftarrow\{(p_{S},o_{i,j},R_{i,j}^{S}):q_{i}\in\mathcal{Q}_{S},1\leq j\leq G\}$ where $R_{i,j}^{S}=\text{score}_{\text{correct}}(o_{i,j},a_{i,j})$
21:  Update $\pi_{\theta}^{(t+1)}\leftarrow\text{GRPO}(\pi_{\theta}^{(t)},\mathcal{D}_{T}\cup\mathcal{D}_{S})$
22:  $\mathcal{P}_{t}\leftarrow\mathcal{P}_{t-1}\cup\mathcal{Q}_{\text{valid}}$
23:end for
24:return $\pi_{\theta}^{(T)}$
@@ 3.1 Training Setup
We experiment with four instruction-tuned models: Llama-3.2-3B-Instruct, Llama-3.1-8B-Instruct  (Dubey et al., 2024), Gemma-2-2B-Instruct (Team et al., 2024), and Qwen-2.5-3B-Instruct (Team, 2024) with GRPO (Shao et al., 2024). We use a learning rate of $3\times 10^{-7}$ and 10 warm-up steps. The KL divergence coefficient is set to $10^{-4}$ and the batch size is 256. To compare models trained on the same number of problem-solution pairs, we train the GRPO baselines with 100 steps, and OpenSIR for 200 steps since OpenSIR allocates half of its training budget to problem generation. Clipping is not applied since we strictly use on-policy samples. Each experiment is run with three random seeds. We provide full training details in Appendix D.1.
@@ 3.3 Baselines
(3) Absolute Zero (Zhao et al., 2025) A self-play framework for code generation that uses Python as external verifier, rewarding problems with minimal solve rates.
(4) R-Zero (Huang et al., 2025) A verifier-free self-play framework that trains separate challenger and solver models using rewards based on repetition penalties and solve rates near 0.5. OpenSIR differs by explicitly optimising for diversity and incorporating solution length to capture multiple dimensions of difficulty within a single model.
@@ 3.4 Main Results
Absolute Zero 84.62 65.33 25.21 48.31 28.12 50.32+0.08
R-Zero 84.22 64.93 24.81 48.45 27.82 50.05-0.19
OpenSIR 85.38 65.87 25.96 48.74 28.33 50.85+0.61
Llama-3.1-8B-Instruct
Base 84.50 47.89 22.75 34.10 16.26 41.10
[…]
Absolute Zero 86.89 51.38 23.21 34.39 15.96 42.37+1.27
R-Zero 86.19 50.93 24.11 32.93 15.66 41.96+0.86
OpenSIR 87.30 52.38 27.29 36.29 17.81 44.21+3.11
Table 1: The avg@16 performance on five mathematical benchmarks. OpenSIR outperforms GRPO baselines trained on >7,000 human-annotated examples and other self-play methods (Absolute Zero, R-Zero) across model families, starting from a single trivial seed problem.
Table 1 demonstrates that OpenSIR achieves substantial gains over the base instruction models across different model scales and families. OpenSIR improves Llama-3.2-3B-Instruct by 3.6 points, Llama-3.1-8B-Instruct by 3.1, and Gemma-2-2B-Instruct by 6.3 points on average accuracy. One exception is Qwen-2.5-3B-Instruct (+.0.6), where all methods show limited gains. The limited improvement aligns with observations of potential benchmark contamination (Wu et al., 2025).
OpenSIR outperforms all GRPO baselines without using human-annotated training data. GRPO baselines require over 7,000 labeled examples, yet OpenSIR generates its own training problems through self-play, starting from a single trivial seed problem. OpenSIR also substantially outperforms other self-play methods by 1.75 to 3.38 points on Llama-3.2-3B-Instruct, Gemma-2-2B-Instruct, and Llama-3.1-8B-Instruct. Although Absolute Zero and R-Zero demonstrate significant improvements over non-instruction-tuned models in their original work (Zhao et al., 2025; Huang et al., 2025), both show limited gains on instruction-tuned models. This challenging scenario is precisely where self-play methods are intended to excel - when models have already consumed available human-annotated data. As we reveal later in our analysis, the success of OpenSIR can be attributed to its ability to explore diverse mathematical concepts, and calibrating difficulty adaptively to maintain optimal challenge levels (§4.1). These capabilities enable OpenSIR to self-improve and expand its skills without external training data, achieving open-ended learning.
@@ A.1 Full Results
Table 5: Math reasoning evaluation results for 2B/3B models with individual seed reporting. We report avg@16 per problem for each seed (42, 43, 44) and their average with standard deviation as superscript.
Models Seed GSM8K MATH-500 Minerva College Olympiad- Avg.
Math Bench
Llama-3.1-8B-Instruct
Base - 84.50 47.89 22.75 34.10 16.26 41.10
[…]
43 88.30 49.93 24.45 34.33 16.29 42.66
44 88.07 50.29 24.43 34.60 17.35 42.95
[…]
43 85.98 50.53 23.88 34.48 16.56 42.29
44 85.77 50.91 23.63 34.57 15.93 42.16
[…]
OpenSIR 42 88.05 52.62 27.79 37.14 18.45 44.81
43 87.03 52.07 27.15 35.82 17.81 43.98
44 86.82 52.45 26.94 35.91 17.18 43.86
[…]
Table 6: Math reasoning evaluation results for 8B models with individual seed reporting. We report avg@16 per problem for each seed (42, 43, 44) and their average with standard deviation as superscript.
We provide the full results of all seeds in Table 5 and 6.
@@ A.3 Further Analysis on Questions Diversity
A.4 Further Analysis on Question Difficulty Progression
Figure 14: Progression of solve rates of OpenSIR and the initial instruction model as training goes.
[…]
A.5 Sensitivity to the initial seed problem
To address whether OpenSIR can robustly escape the limited starting point of a trivial arithmetic seed (“What is 1+1?”), we experiment with two substantially different initial seeds using Llama-3.2-3B-Instruct: a geometry problem from the MATH dataset, representing a different mathematical domain, and a competition-level problem from AIME 2024, which is significantly more challenging than the trivial seed.
$B$ and $C$ trisect $\overline{AD}$ and $M$ is the midpoint of $\overline{AD}$. $MC=8$. How many units are in the length of $\overline{AD}$?
Figure 15: A geometry problem from the MATH dataset, representing a different mathematical domain from the trivial arithmetic seed.
Every morning Aya goes for a 99-kilometer-long walk and stops at a coffee shop afterwards. When she walks at a constant speed of $s$ kilometers per hour, the walk takes her 4 hours, including $t$ minutes spent in the coffee shop. When she walks $s+2$ kilometers per hour, the walk takes her 2 hours and 24 minutes, including $t$ minutes spent in the coffee shop. Suppose Aya walks at $s+\frac{1}{2}$ kilometers per hour. Find the number of minutes the walk takes her, including the $t$ minutes spent in the coffee shop.
Figure 16: A competition-level problem from AIME 2024, significantly more challenging than the trivial seed.
Model Acc
OpenSIR 38.42
[…]
Table 7: Performance of OpenSIR with different initial seed problem.
Table 7 shows that all three variants achieve nearly identical performance (38.42, 38.67, and 38.81), with differences of less than 0.5 percentage points. This demonstrates that OpenSIR is robust to the initial seed problem, successfully escaping the limited starting point regardless of whether it begins with trivial arithmetic, a different mathematical domain (geometry), or a significantly more challenging competition-level problem.
Figure 17: t-SNE visualisation of problem embeddings generated by OpenSIR from three different initial seeds. The substantial overlap demonstrates that the method converges to similar problem distributions regardless of the starting point.
Figure 17 visualises the diversity of problems generated at the final training step across the three different initial seeds. The t-SNE embeddings reveal that all three variants produce diverse problems spanning similar regions of the semantic space, with substantial overlap in their distributions regardless of the initial seed. This confirms that OpenSIR successfully escapes its starting point by exploring a wide range of mathematical concepts, driven by the diversity and solvability rewards that encourage continuous exploration beyond the initial problem domain and difficulty level.
A.6 OpenSIR Incentivises Reasoning Capacity
To verify whether OpenSIR elicits genuine reasoning improvements rather than memorisation, we evaluate pass@k performance on five challenging mathematical benchmarks following (Yue et al., 2025).
Figure 18 shows that OpenSIR consistently outperforms base instruction models across all k values (8–256) on all benchmarks. These results confirm that OpenSIR drives genuine advances in mathematical reasoning capacity.
Figure 18: Pass@k curves comparing base instruction models and OpenSIR across five mathematical benchmarks. OpenSIR consistently improves performance across all k values, with stable or increasing gaps at higher k, demonstrating genuine reasoning improvements rather than memorization.
A.7 Prolonged Training Analysis
Figure 19: Performance of OpenSIR extended training using Llama-3.2-3B-Instruct.
To better understand the limitation of OpenSIR, we extend training of a single run of Llama-3.2-3B-Instruct to 500 steps, substantially beyond the 200 steps used in the main experiments. Figure 19 shows the evaluation performances over training. We observe consistent improvement from 34.8% to 41.3% at step 350, representing a gain of +6.5 points. Performance then plateaus after step 350 and remains stable around 40-41% through step 500, with no further statistically significant gains.
Preliminary examination of the generated problems suggests a likely cause for this saturation. The model appears to have explored most major mathematical topics by step 350, after which the generated problems become increasingly similar and repetitive. This indicates that the diversity reward mechanism may become less effective over extended training. Future work could investigate more sophisticated diversity mechanisms to foster open-ended exploration over long training horizons.
A.8 Synergy with Annotated Data
Model GSM8K MATH-500 Minerva College OlympiadBench Avg.
Math
Base 73.94 42.86 15.21 28.78 13.09 34.78
[…]
OpenSIR 78.28 46.22 17.46 34.42 15.72 38.42+3.64
GSM8K $\rightarrow$ OpenSIR 81.43 46.12 19.43 36.15 18.35 40.30+5.52
GSM8K & OpenSIR 81.57 49.48 20.39 36.85 18.14 41.29+6.51
Table 8: The avg@16 performance on five mathematical benchmarks. OpenSIR obtains better results when trained together with GSM8K compared to OpenSIR or GSM8K alone.
While we showed that OpenSIR achieves significant improvements in math reasoning without using annotated data, we further investigate if OpenSIR can be combined with annotated data to achieve even greater performance gains. Having demonstrated that OpenSIR achieves significant improvements without annotated data, we investigate whether combining OpenSIR with annotated data can yield further gains. We focus on Llama-3.2-3B-Instruct and use Gsm8K as the training data, as Table 1 shows that fine-tuning on GSM8K consistently outperforms using MATH.
We explore two training strategies: (1) GSM8K $\rightarrow$ OpenSIR: The model is first trained on GSM8K for half the training iterations, then trained with OpenSIR for the remaining half. (2) GSM8K & OpenSIR: Each training iteration uses half GSM8k samples and half OpenSIR samples.
Table 8 shows that both setups achieve better performance than using OpenSIR alone or only on GSM8K. One possible explanation for the sequential approach’s effectiveness (GSM8K $\rightarrow$ OpenSIR) is that training on GSM8K first may improve the model’s foundational reasoning abilities, which could provide a stronger starting point for OpenSIR’s self-generated questions. The concurrent approach (GSM8K & OpenSIR) achieves a slight additional edge, which might be attributed to the model receiving better feedback signals for question calibration from the beginning, as it can leverage both supervised and self-generated data simultaneously throughout training. The precise underlying mechanisms for these improvements require further investigation.
A.9 Computational Cost Analysis
In standard GRPO training, each iteration processes a batch of $B$ prompts, generating $G$ responses for each prompt, resulting in $B\times G$ total forward passes per iteration. In method, each training iteration involves generating $B$ problems and $G$ solution attempts for each problem, yielding $B$ forward passes for problem generation and $B\times G$ forward passes for solution generation, for a total of $B+B\times G=B(1+G)$ forward passes. Compared to the $B\times G$ forward passes in standard GRPO training, method requires an additional $B$ forward passes for problem generation. This represents a relative computational overhead of $\frac{B}{B\times G}=\frac{1}{G}$, or 12.5% with $G=8$ solution attempts per problem.
Problem embeddings for diversity scoring are computed asynchronously during solution generation, incurring no additional wall-clock time. Cosine distance calculations between problem embeddings require $\mathcal{O}(B\times|\mathcal{P}_{t}|)$ operations where $|\mathcal{P}_{t}|$ is the problem pool size, but execute in under 3 seconds per iteration in our experiments—negligible compared to LLM forward passes. Note that while method generates $B$ problems per iteration, only the top-scoring $B/2$ problems are selected for teacher training and another top-scoring $B/2$ for student training, as described in Section 2.4. Overall, method achieves improved performance without human-annotated training data at this modest computational overhead.